OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x a)2 sin x cos2x + cos x b)2 sin x cos2x + sin3x c)sin x cos2x - sin3x + cos3x d)2 cos2x sin x + sin x - 2 sin3x

OpenStudy (amoodarya):

sin(3x)=sin(x+2x) =sinx cos 2x + cos x sin 2x =sinx (2cos^2x -1)_cosx (2sinx cos x)

OpenStudy (anonymous):

so its c?

OpenStudy (amoodarya):

whats the original question?

OpenStudy (anonymous):

the original question is Rewrite with only sin x and cos x. sin 3x

OpenStudy (amoodarya):

ok you mean write sin3x by only (sinx , cos x) ?

OpenStudy (anonymous):

yes

OpenStudy (amoodarya):

then whats mean a,b,c,d because they use sin 2x or sin3x ?

OpenStudy (anonymous):

thos are the answer choice

OpenStudy (amoodarya):

c is not correct because if you put x=0 at sin3x=0 c(0)=1 a(0)=1

OpenStudy (anonymous):

oh ok so what is it?

OpenStudy (amoodarya):

let f(x)=sin 3x answer is equal to f(x) ok so for example f(0)=0 or f(pi/2)=sin(3pi/2)=-1 if you put pi/2 in them only (d) is answer

OpenStudy (amoodarya):

understand ?

OpenStudy (anonymous):

yes thankyou!

OpenStudy (amoodarya):

i think that ques. has a logic problem i was teaching trigonometry for many years